# Defending Against Sophisticated Poisoning Attacks with RL-based Aggregation in Federated Learning

### v1

@@ Abstract
Federated learning is highly susceptible to model poisoning attacks, especially those meticulously crafted for servers. Traditional defense methods mainly focus on updating assessments or robust aggregation against manually crafted myopic attacks. When facing advanced attacks, their defense stability is notably insufficient. Therefore, it is imperative to develop adaptive defenses against such advanced poisoning attacks. We find that benign clients exhibit significantly higher data distribution stability than malicious clients in federated learning in both CV and NLP tasks. Therefore, the malicious clients can be recognized by observing the stability of their data distribution. In this paper, we propose AdaAggRL, an RL-based Adaptive Aggregation method, to defend against sophisticated poisoning attacks. Specifically, we first utilize distribution learning to simulate the clients’ data distributions. Then, we use the maximum mean discrepancy (MMD) to calculate the pairwise similarity of the current local model data distribution, its historical data distribution, and global model data distribution. Finally, we use policy learning to adaptively determine the aggregation weights based on the above similarities. Experiments on four real-world datasets demonstrate that the proposed defense model significantly outperforms widely adopted defense models for sophisticated attacks.
1 Introduction
@@
Federated Learning (FL) is a promising machine learning method where model training is distributed across multiple local devices rather than centralized on a single server. It can not only ensure the security of users’ private data but also enhance model performance by collaborating diverse data from various local sources. It finds applications in diverse scenarios, including smart healthcare Chaddad et al. (2023), financial services Byrd and Polychroniadou (2020), IoT Nguyen et al. (2021), and intelligent transportation Yamany et al. (2021). However, FL systems are vulnerable, and the performance of the aggregated model is highly susceptible to model poisoning attacks from unknown clients, especially the sophisticated poisoning strategies tailored for central servers. In this work, we focus on untargeted model poisoning attacks, where malicious devices aim to maximally reduce the accuracy of the global model by sending customized gradients to the server.
@@
Traditional defense methods mainly rely on designing local model update assessment mechanisms or using robust aggregation methods to mitigate the impact of poisoning attacks. However, these defense methods are primarily targeted at manually crafted myopic attack strategies, and their defense stability is noticeably lacking when facing advanced attacks. For example, Li et al. (2022) propose using distribution learning to simulate the data distribution of the central server and employing reinforcement learning (RL) to tailor attack policy for the aggregation process, making it less detectable, as shown in Figure 5. Therefore, it is urgent to develop adaptive defenses against such learnable advanced poisoning attacks.
@@
Most benign clients exhibit significant data distribution stability in federated learning. In normal cases, the current training data distribution simulated by the parameters passed through the client should align with its historical simulated data distribution. This is because benign clients typically employ random sampling and undergo multi-round training based on their local data, ensuring the stability of the training data distribution used for global updates. However, malicious clients require gradient attacks, so its simulated data distribution between current and history lacks regularity. To validate this, we conduct a statistical analysis of mainstream attack methods, namely IPM Xie et al. (2020), LMP Fang et al. (2020), EB Bhagoji et al. (2019) and RL-based attacks Li et al. (2022), measuring the similarity of their data distribution with history after each round of updates on MNIST dataset, as shown in Figure LABEL:fig:statistic. We find that benign clients maintain higher data distribution similarity, while attack clients show no discernible patterns. We also observe the same phenomenon on NLP tasks, as shown in Appendix E. Therefore, malicious clients can be recognized by observing the stability of their simulated data distribution.
@@
In this paper, we propose an RL-based Adaptive Aggregation method, AdaAggRL, to thwart sophisticated poisoning attacks. It determines the aggregation weights of local models by comparing the stability of client data distributions. Specifically, we first use distribution learning to emulate client data distributions based on the uploaded model parameters. Next, we use the maximum mean discrepancy (MMD) to calculate the pairwise similarity of the current local model data distribution, its historical data distribution, and the global model data distribution, to evaluate the stability of the client’s data distribution. Considering that the accuracy of distribution learning can potentially impact the calculation of the similarities above, we use reconstruction similarity as an evaluation metric for the quality of distribution learning. Finally, we use the policy learning method TD3 to adaptively determine the aggregation weights based on the above three distribution similarities and the reconstruction similarity of distribution learning.
@@
Experimental results on four real-world datasets demonstrate that the proposed AdaAggRL defense method significantly outperforms existing defense approaches Blanchard et al. (2017); Yin et al. (2018); Sun et al. (2019) and achieves a more stable global model accuracy even facing sophisticated attacks, such as RL-based attacks Li et al. (2022).
@@
  • We propose an RL-based adaptive aggregation method AdaAggRL to defend against sophisticated untargeted model poisoning attacks tailored for servers in federated learning, aiming to advance the development of defense systems.
@@
2 Motivation
@@
In most cases, benign clients’ data distribution remains stable, while malicious clients lack regularity. During FL, benign clients are selected by the server through random sampling and trained on their local data for a certain number of rounds. Therefore, the simulated data distribution obtained through gradient inversion should align with their historical distribution. But the data distribution of malicious clients lacks regularity, due to their need to conduct model attacks. To verify this, we conduct a statistical analysis of the similarity between data distributions for mainstream attack methods such as IPM, LMP, EB, and RL-attack, as shown in Figure LABEL:fig:statistic. The similarity of data distributions for benign clients remains high and stable, while malicious clients lack any regular pattern. Though the stability of advanced attack RL-attack is significantly higher than other ordinary attack methods, there are still significant differences compared to benign clients. We also observe the same phenomenon on NLP tasks, as shown in Appendix E.
@@
Moreover, we observe a certain degree of similarity between current client data distribution and global model data distribution, as shown in Figure LABEL:fig:2. The data distribution of benign clients for current steps is similar to the global model and remains stable, but the similarity of malicious clients is lower. Since the data distribution of the global model represents the average state of normal data, the distribution of benign clients should always be consistent with the global model. Instead, malicious clients do not possess this property. Though the advanced RL-attack has high similarity with the global model due to collecting the data distribution of the server, it is still not as stable as benign clients. Therefore, we can compare data distribution similarity between the current client and global model to observe variations and assess the degree of malicious behavior. Similarly, the similarity between historical distribution and global model distribution is higher and more stable for benign clients than malicious ones, as shown in Appendix C.
@@
Considering that the quality of distribution learning greatly affects the accurate calculation of these three distribution similarities, the reconstruction similarity of distribution learning is used as an evaluation metric for assessing the quality of distribution learning. Reconstruction similarity measures the similarity between the inversion gradients obtained from distribution learning and the gradients updated by clients. Higher reconstruction similarity indicates a higher level of confidence in current distribution learning.
3 RL-based Adaptive Aggregation Methods
3.1 Task Definition
@@
In the context of FL McMahan et al. (2017), a system comprises $K$ clients, with each client $k$ possessing a fixed local dataset $D_{k}=\{(x_{kj},y_{kj})\}_{j=1}^{n_{k}}$, where $n_{k}$ is the size of $D_{k}$. The local objective of client $k$ is $F_{k}(\theta)=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}l(\theta,x_{kj},y_{kj})$, where $l$ is the loss function. And $(x_{kj},y_{kj})$ is the $j$-th data sample drawn i.i.d. from some distribution $P_{k}$. $\hat{P}_{k}$ denotes the empirical distribution of $n_{k}$ data samples. The optimization objective of FL is: $\min_{\theta}f(\theta)=\sum_{k=1}^{K}p_{k}F_{k}(\theta)$, $p_{k}$ represents the weight of client $k$. During FL, in each epoch $t\geq 0$, the server randomly selects a subset $C^{t}$ from all clients and distributes the latest global model parameters $\theta^{t}$. The chosen clients train on their local datasets, updating parameters as $\theta^{t+1}_{k}=\theta^{t}-\alpha\nabla F_{k}(\theta)$, where $\alpha$ is the learning rate. Then they upload the updated model parameters. The server aggregates the received parameters according to a specific aggregation rule, denoted as $Aggr$, to obtain the new global model parameters $\theta^{t+1}=Aggr(\theta^{t+1}_{C^{t}})$.
@@
3.2 Framework Overview
Figure 4: An overview of RL-attack
@@ 3.2 Framework Overview
In AdaAggRL framework (see Algorithm in Appendix A), the server determines the weights for local model aggregation by assessing the stability of client data distributions, as shown in Figure 5. Firstly, we employ distribution learning to simulate the client’s data distribution $\hat{P}^{t}_{k}$ based on the locally uploaded model parameters $\theta^{t+1}_{k}$. Secondly, we calculate similarity metrics $S_{k,cl}$ between the current client and historical data distribution, $S_{k,cg}$ between the current client and global model data distribution, and $S_{k,lg}$ between the client’s historical and global model data distribution. Finally, RL is utilized to adaptively determine the weight $p_{k}$ for local model aggregation based on these metrics and the reconstruction similarity $S_{k,R}$ from distribution learning.
3.3 Distribution Learning
@@
According to local model parameters $\theta^{t+1}_{k}$ uploaded by clients, the server simulates the local data distribution $\hat{P}^{t}_{k}$ of clients by gradient inversion. In this work, we adapt the inverting gradients (IG) method Geiping et al. (2020) for distribution learning. The IG method reconstructs the data sample by optimizing the loss function based on the cosine similarity between the real gradient and the gradient generated by the reconstructed data.
@@
For each epoch $t\geq 0$, the server receives the clients’ locally updated model parameters and calculates the corresponding batch level gradient $\bar{g}^{t}_{k}\coloneqq(\theta^{t+1}_{k}-\theta^{t})/\alpha$. The server then solves the following optimization problem with a batch of randomly generated dummy data and labels $D_{dummy}$: $\min_{D_{\text{dummy}}}1-\frac{\langle\nabla_{\theta}F_{D_{\text{dummy}}}(\theta^{t+1}_{k}),\bar{g}^{t}_{k}\rangle}{\|\nabla_{\theta}F_{D_{\text{dummy}}}(\theta^{t+1}_{k})\|\cdot\|\bar{g}^{t}_{k}\|}+\frac{\beta}{B^{\prime}}\Sigma_{(x,y)\in D_{\text{dummy}}}\mathrm{TV}(x)$, where $\beta$ is a fixed parameter, $B^{\prime}$ is the size of the dummy data batch, $F_{D_{\text{dummy}}}(\theta)=\frac{1}{B^{\prime}}\sum_{(x,y)\in D_{\text{dummy}}}l(\theta,x,y)$, $\mathrm{TV}$ calculates the total variationRudin et al. (1992). While solving the optimization problem, $D_{dummy}$ is continuously updated. The optimization terminates after $max\_iters$ iterations, then outputs the updated data as the reconstructed data samples $D_{\text{rec}}$, and the reconstruction similarity of client $k$, denoted as $S_{k,R}=\frac{\langle\nabla_{\theta}F_{D_{\text{rec}}}(\theta^{t+1}_{k}),\bar{g}^{t}_{k}\rangle}{\|\nabla_{\theta}F_{D_{\text{rec}}}(\theta^{t+1}_{k})\|\cdot\|\bar{g}^{t}_{k}\|}$.
3.4 Environmental Cues
@@
The distribution of samples is extracted by employing a pre-trained CNN to convert the image samples $D_{\text{rec}}$ into a collection of feature vectors $V$. For each client $k$, the difference $d_{k,cl}$ between this feature distribution and the history distribution is calculated using the maximum mean discrepancy (MMD) Arbel et al. (2019); Wang et al. (2021) as $d_{k,cl}=\mathrm{MMD}(V^{\text{current}}_{k},V^{\text{history}}_{k})\in[0,+\infty)$. Here, $V^{\text{current}}_{k}$ and $V^{\text{history}}_{k}$ are feature vectors of current and historical data distributions respectively. $V_{g}$ represents feature vectors of the global model data distribution, obtained by averaging feature vectors from all participating clients. Subsequently, the dissimilarity between the current data distribution of client $k$ and the global one is $d_{k,cg}=\mathrm{MMD}(V^{\text{current}}_{k},V_{g})$. Similarly, the dissimilarity between the historical data distribution and the global one is $d_{k,lg}=\mathrm{MMD}(V^{\text{history}}_{k},V_{g})$. And the similarity between the current data distribution and the historical data distribution $S_{k,cl}$ is calculated using the following formula:
@@
3.5 Actions Learning
The server dynamically adapts the aggregation weights based on three obtained similarity metrics and the reconstruction similarity associated with each client. By utilizing experiences sampled from the simulated environment, the server engages in learning a collaborative defense strategy aimed at minimizing empirical loss. This learning process employs the RL algorithm TD3 Fujimoto et al. (2018).
@@ 3.5 Actions Learning
3.6 Computational Complexity
Compared to classical FL, AdaAggRL’s increased computational complexity mainly stems from the $\mathrm{EnvironmentalCues}$ step. In IG algorithm, computation involves model forward propagation, loss function calculation, and backpropagation for optimization. The computational complexity depends on model complexity $M$, optimization rounds $max\_iters$, and reconstructed images $num\_images$, totaling $O(max\_iters\times M\times num\_images)$. Extracting image features via pre-trained CNN incurs a complexity of $O(C_{CNN}\times num\_images)$. Considering MMD’s constant complexity $O(C_{MMD})$, the overall computational complexity for $\mathrm{EnvironmentalCues}$ is $O(|C^{t}|\times num\_images\cdot(M\times max\_iters+C_{CNN}))+O(3|C^{t}|\cdot C_{MMD})+O(1)$.
@@ 3.6 Computational Complexity
AdaAggRL Krum Median C-Median FLtrust Clipping
MNIST 1.6520 1.7438 1.0690 1.5556 1.1066 1.3265
Cifar10 7.2009 8.6983 3.5374 4.4695 3.5892 3.6956
@@
4 Experiments
4.1 Experimental Settings
4.1.1 Dataset
@@
We conduct experiments on four datasets: MNIST LeCun et al. (1998), F-MNIST Xiao et al. (2017), EMNIST Cohen et al. (2017), and Cifar10 Krizhevsky et al. (2009). Addressing the non-i.i.d. challenge in FL, we follow the approach from prior work Fang et al. (2020) by distributing training examples across all clients. Given an M-class dataset, clients are randomly divided into M groups. The probability $q$ of assigning a training sample with label $l$ to its respective group is set, with the probability of assigning it to other groups being $\frac{1-q}{M-1}$. Training samples within the same client group adhere to the same distribution. When $q=1/M$, the distribution of training samples across M groups is uniform, ensuring that all clients’ datasets follow the same distribution. In cases where $q>1/M$, the datasets among clients are not identically distributed. Using MNIST dataset, we set $q=0.5$ to distribute training samples among clients unevenly, denoted as MNIST-0.5. MNIST-0.1 represents the scenario where MNIST is evenly distributed among clients ($q=0.1$).
4.1.2 Metrics
@@
We assess FL defense methods by evaluating the global model’s image classification accuracy after 500 epochs of FL training since these attacks aim to diminish testing accuracy. The classification accuracy is the ratio of correctly predicted test examples to the total number of test examples. Higher testing accuracy of the global model under various attacks indicates stronger robustness of the defense.
4.1.3 Baselines
@@
We compare AdaAggRL with five other defense algorithms: Krum Blanchard et al. (2017), coordinate-wise median (Median) Yin et al. (2018), norm clipping (Clipping) Sun et al. (2019), an extension of the vanilla coordinate-wise median (C-Median) where a norm clipping step is applied Li et al. (2022), and FLtrust Cao et al. (2020). Krum filters malicious updates at the client level, Clipping performs gradient clipping on parameter updates before aggregation, Median and C-Median select the median or clipped median of individual parameter values from all local model updates as global model parameters, and FLtrust requires the server to have access to a small amount of root data.

### v2

@@ Abstract
Federated learning is susceptible to model poisoning attacks, especially those meticulously crafted for servers. Traditional defense methods mainly focus on updating assessments or robust aggregation against manually crafted myopic attacks. When facing advanced attacks, their defense stability is notably insufficient. Therefore, it is imperative to develop adaptive defenses against such advanced poisoning attacks. We find that benign clients exhibit significantly higher data distribution stability than malicious clients in federated learning in both CV and NLP tasks. Therefore, the malicious clients can be recognized by observing the stability of their data distribution. In this paper, we propose AdaAggRL, an RL-based Adaptive Aggregation method, to defend against sophisticated poisoning attacks. Specifically, we first utilize distribution learning to simulate the clients’ data distributions. Then, we use maximum mean discrepancy (MMD) to calculate the pairwise similarity of the current local model data distribution, its historical data distribution, and global model data distribution. Finally, we use policy learning to adaptively determine the aggregation weights based on the above similarities. Experiments on four real-world datasets demonstrate that the proposed defense model significantly outperforms widely adopted defense models for sophisticated attacks.
Introduction
@@
Federated Learning (FL) enables distributed model training across local devices, preserving data privacy while leveraging diverse local data to improve performance. It is widely applied in areas including smart healthcare (Chaddad, Wu, and Desrosiers 2023), financial services (Byrd and Polychroniadou 2020), IoT (Nguyen et al. 2021), and intelligent transportation (Yamany, Moustafa, and Turnbull 2021). However, FL systems are vulnerable, and the performance of the aggregated model is susceptible to model poisoning attacks from unknown clients (Zheng et al. 2024), especially the sophisticated poisoning strategies tailored for central servers. In this work, we focus on untargeted model poisoning attacks, where malicious devices aim to maximally reduce the accuracy of the global model by sending customized gradients to the server.
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 1: The statistical results of the similarity between the current client data distribution and its historical data distributions under four types of attacks vary with the training epochs on MNIST dataset. The x-axis denotes the number of client update rounds, and the y-axis represents the similarity between the current and its historical data distributions.
@@
Traditional defense methods mainly rely on designing local model update assessment mechanisms or using robust aggregation methods to mitigate the impact of poisoning attacks. However, these defense methods are primarily targeted at manually crafted myopic attack strategies, and their defense stability is lacking when facing advanced attacks. For example, Li, Sun, and Zheng propose using distribution learning to simulate the data distribution of the central server and employing reinforcement learning (RL) to tailor attack policy for the aggregation process, making it less detectable.Therefore, it is urgent to develop adaptive defenses against such learnable advanced poisoning attacks.
@@
Most benign clients exhibit significant data distribution stability in FL. In normal cases, the current training data distribution simulated by the client’s parameters should align with its historical simulated data distribution. This is because benign clients typically employ random sampling and undergo multi-round training based on their local data, ensuring the stability of the simulated data distribution. However, malicious clients require gradient attacks, so its simulated data distribution between current and history lacks regularity. To validate this, we conduct a statistical analysis of mainstream attack methods, namely IPM (Xie, Koyejo, and Gupta 2020), LMP (Fang et al. 2020), EB (Bhagoji et al. 2019) and RL-based attacks (Li, Sun, and Zheng 2022), measuring the similarity of their data distribution with history after each round of updates on MNIST, as shown in Figure 1. We find that benign clients maintain higher data distribution similarity, while attack clients show no discernible patterns. We also observe the same phenomenon on NLP tasks, as shown in Appendix. Therefore, malicious clients can be recognized through the stability of their simulated data distribution.
@@
This paper proposes an RL-based Adaptive Aggregation method, AdaAggRL, to thwart sophisticated poisoning attacks. It determines aggregation weights of local models by comparing the stability of client data distributions. Specifically, we first use distribution learning to emulate client data distributions based on the uploaded model parameters. Next, we use the maximum mean discrepancy (MMD) to calculate the pairwise similarity of the current local model data distribution, its historical data distribution, and the global model data distribution, to evaluate the stability of the client’s data distribution. Considering that the accuracy of distribution learning can potentially impact the calculation of the similarities above, we use reconstruction similarity as an evaluation metric for the quality of distribution learning. Finally, we use the policy learning method TD3 to adaptively determine the aggregation weights based on these similarities.
@@
Experimental results on four real-world datasets demonstrate that AdaAggRL defense method significantly outperforms existing defense approaches (Blanchard et al. 2017; Yin et al. 2018; Sun et al. 2019) and achieves a more stable global model accuracy even facing sophisticated attacks, such as RL-based attacks (Li, Sun, and Zheng 2022).
@@
  • We propose an RL-based adaptive aggregation method AdaAggRL to defend against sophisticated untargeted model poisoning attacks tailored for servers in FL, aiming to advance the development of defense systems.
@@
Related Work
Poisoning Attacks
Based on the attacker’s objectives, poisoning attacks can be classified into targeted poisoning attacks aiming to misclassify specific input sets (Bhagoji et al. 2019; Baruch, Baruch, and Goldberg 2019; Bagdasaryan et al. 2020) and untargeted attacks aimed at reducing the overall accuracy of the global model (Fang et al. 2020; Xie, Koyejo, and Gupta 2020; Shejwalkar and Houmansadr 2021). Current untargeted attack methods typically employ heuristic-based approaches (Xie, Koyejo, and Gupta 2020) or optimize myopic objectives (Fang et al. 2020; Shejwalkar and Houmansadr 2021; Shejwalkar et al. 2022). Bhagoji et al. generate malicious updates through explicit enhancement, optimizing for a malicious objective strategically designed to induce targeted misclassification. Xie, Koyejo, and Gupta manipulate the attacker’s gradients to ensure the inner product with the true gradients becomes negative. Fang et al. generate malicious updates by solving an optimization problem. However, these attack methods require local updates from benign clients or accurate global model parameters to generate significant adversarial impacts and often yield suboptimal results when robust aggregation rules are employed.
To address these deficiencies, Li, Sun, and Zheng propose a model-based RL framework for guiding untargeted poisoning attacks in FL system. They utilize the server’s updates to approximate the server’s data distribution, subsequently employing the learned distribution as a simulator for FL environment. Based on the simulator, they utilize RL to automatically generate effective attacks, resulting in significantly reducing the global accuracy. Even when the server employs robust aggregation rules, this customized method can still maintain a high level of attack effectiveness.
Defenses for Poisoning Attacks
Current defense strategies against FL poisoning attacks can be categorized into two types: one involves designing local model update evaluation mechanisms to identify malicious client-submitted model parameters (Cao et al. 2020; Sattler et al. 2020; Zhang et al. 2022), and the other is based on designing novel Byzantine fault-tolerant aggregation algorithms using mathematical statistics to improve the robustness of aggregation (Blanchard et al. 2017; Yin et al. 2018; Sun et al. 2019; Rajput et al. 2019; Xie, Koyejo, and Gupta 2019). In evaluation mechanisms, Sattler et al. divide updates into different groups based on cosine similarity between the model parameters submitted by clients, mitigating the impact of poisoning attacks. In response to more covert poisoning attacks, some evaluation methods (Cao et al. 2020; Park et al. 2021) require the server to collect a portion of clean data as a basis for validating model updates. In robust aggregation algorithms, statistical methods (Blanchard et al. 2017; Yin et al. 2018; Sun et al. 2019) compare local updates and remove statistical outliers before updating the global model. For example, Blanchard et al. employ the square-distance metric to measure distances among local updates and select the local update with the minimum distance as the global parameters. Yin et al. sort the values of parameters in all local model updates and consider the median value of each parameter as the corresponding parameter value in the global model update. Sun et al. perform gradient clipping on parameter updates before aggregation. Due to the susceptibility of statistical estimates to an outlier, existing aggregation methods cannot guarantee accuracy well and are still susceptible to local model poisoning attacks (Bhagoji et al. 2019; Fang et al. 2020).
Motivation
@@
During FL, benign clients are selected by the server through random sampling and trained on their local data for a certain number of rounds. Therefore, the simulated data distribution obtained through gradient inversion should align with their historical distribution. But the data distribution of malicious clients lacks regularity, due to their need to conduct model attacks. To verify this, we conduct a statistical analysis of the similarity between data distributions for mainstream attack methods such as IPM, LMP, EB, and RL-attack, as shown in Figure 1. The similarity of data distributions for benign clients remains high and stable, while malicious clients lack any regular pattern. We also observe the same phenomenon on NLP tasks, as shown in Appendix.
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 2: The statistical results of the similarity between the current client data distribution and the global model data distribution under attacks vary with epochs on MNIST.
@@
Moreover, the data distribution of benign clients for current steps is similar to the global model and remains stable, but the similarity of malicious clients is lower, as shown in Figure 2. Since the data distribution of the global model represents the average state of normal data, the distribution of benign clients should always be consistent with the global model. Instead, malicious clients do not possess this property. Therefore, we can compare data distribution similarity between the current client and global model to observe variations and assess the degree of malicious behavior. Similarly, the similarity between historical distribution and global model distribution is higher and more stable for benign clients than malicious ones, as shown in Appendix.
@@
Since the quality of distribution learning greatly affects the accuracy of distribution similarities, the reconstruction similarity of distribution learning is used as an evaluation metric for assessing the quality of distribution learning. Reconstruction similarity measures the similarity between the inversion gradients from distribution learning and the gradients updated by clients. Higher reconstruction similarity indicates higher confidence in current distribution learning.
RL-based Adaptive Aggregation Methods
Task Definition
@@
In the context of FL (McMahan et al. 2017), a system comprises $K$ clients, with each client $k$ possessing a fixed local dataset $D_{k}=\{(x_{kj},y_{kj})\}_{j=1}^{n_{k}}$, where $n_{k}$ is the size of $D_{k}$. The local objective of client $k$ is $F_{k}(\theta)=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}l(\theta,x_{kj},y_{kj})$, where $l$ is the loss function. And $(x_{kj},y_{kj})$ is the $j$-th sample drawn i.i.d. from some distribution $P_{k}$. $\hat{P}_{k}$ denotes the empirical distribution of $n_{k}$ data samples. The optimization objective of FL is: $\min_{\theta}f(\theta)=\sum_{k=1}^{K}p_{k}F_{k}(\theta)$, $p_{k}$ represents the weight of client $k$. During FL, in each epoch $t\geq 0$, the server randomly selects a subset $C^{t}$ from all clients and distributes the latest global model parameters $\theta^{t}$. The chosen clients train on their local datasets, updating parameters as $\theta^{t+1}_{k}=\theta^{t}-\alpha\nabla F_{k}(\theta)$, where $\alpha$ is the learning rate. Then they upload the updated model parameters. The server aggregates the received parameters according to a specific aggregation rule $Aggr$, to obtain the new global model parameters $\theta^{t+1}=Aggr(\theta^{t+1}_{C^{t}})$.
@@
Framework Overview
@@
In AdaAggRL framework (see Algorithm in Appendix), the server determines the weights for local model aggregation by assessing the stability of client data distributions, as shown in Figure 3. Firstly, we employ distribution learning to simulate the client’s data distribution $\hat{P}^{t}_{k}$ based on the locally uploaded model parameters $\theta^{t+1}_{k}$. Secondly, we calculate similarity metrics $S_{k,cl}$ between the current client and historical data distribution, $S_{k,cg}$ between the current client and global model data distribution, and $S_{k,lg}$ between the client’s historical and global model data distribution. Finally, RL is utilized to adaptively determine the weight $p_{k}$ for local model aggregation based on these metrics and the reconstruction similarity $S_{k,R}$ from distribution learning.
Distribution Learning
@@
According to local model parameters $\theta^{t+1}_{k}$ uploaded by clients, the server simulates the local data distribution $\hat{P}^{t}_{k}$ of clients by gradient inversion. In this work, we adapt the inverting gradients (IG) method (Geiping et al. 2020) for distribution learning. The IG method reconstructs the data sample by optimizing the loss function based on the cosine similarity between the real gradient and the gradient generated by the reconstructed data.
@@
For each epoch $t\geq 0$, the server receives the clients’ locally updated model parameters and calculates the corresponding batch level gradient $\bar{g}^{t}_{k}\coloneqq(\theta^{t+1}_{k}-\theta^{t})/\alpha$. The server then solves the following optimization problem with a batch of randomly generated dummy data and labels $D_{dummy}$: $\min_{D_{\text{dummy}}}1-\frac{\langle\nabla_{\theta}F_{D_{\text{dummy}}}(\theta^{t+1}_{k}),\bar{g}^{t}_{k}\rangle}{\|\nabla_{\theta}F_{D_{\text{dummy}}}(\theta^{t+1}_{k})\|\cdot\|\bar{g}^{t}_{k}\|}+\frac{\beta}{B^{\prime}}\Sigma_{(x,y)\in D_{\text{dummy}}}\mathrm{TV}(x)$, where $\beta$ is a fixed parameter, $B^{\prime}$ is the size of the dummy data batch, $F_{D_{\text{dummy}}}(\theta)=\frac{1}{B^{\prime}}\sum_{(x,y)\in D_{\text{dummy}}}l(\theta,x,y)$, $\mathrm{TV}$ calculates the total variation(Rudin, Osher, and Fatemi 1992). While solving the optimization problem, $D_{dummy}$ is continuously updated. The optimization terminates after $max\_iters$ iterations, then outputs the updated data as the reconstructed data samples $D_{\text{rec}}$, and the reconstruction similarity of client $k$, denoted as $S_{k,R}=\frac{\langle\nabla_{\theta}F_{D_{\text{rec}}}(\theta^{t+1}_{k}),\bar{g}^{t}_{k}\rangle}{\|\nabla_{\theta}F_{D_{\text{rec}}}(\theta^{t+1}_{k})\|\cdot\|\bar{g}^{t}_{k}\|}$.
Environmental Cues
@@
The distribution of samples is extracted by employing a pre-trained CNN to convert the image samples $D_{\text{rec}}$ into a collection of feature vectors $V$. For each client $k$, the difference $d_{k,cl}$ between this feature distribution and the history distribution is calculated using the maximum mean discrepancy (MMD) (Arbel et al. 2019; Wang et al. 2021) as $d_{k,cl}=\mathrm{MMD}(V^{\text{current}}_{k},V^{\text{history}}_{k})\in[0,+\infty)$. Here, $V^{\text{current}}_{k}$ and $V^{\text{history}}_{k}$ are feature vectors of current and historical data distributions respectively. $V_{g}$ represents feature vectors of the global model data distribution, obtained by averaging feature vectors from all participating clients. Subsequently, the dissimilarity between the current data distribution of client $k$ and the global one is $d_{k,cg}=\mathrm{MMD}(V^{\text{current}}_{k},V_{g})$. Similarly, the dissimilarity between the historical data distribution and the global one is $d_{k,lg}=\mathrm{MMD}(V^{\text{history}}_{k},V_{g})$. And the similarity between the current data distribution and the historical data distribution $S_{k,cl}$ is calculated using the following formula:
@@
Actions Learning
The server dynamically adapts the aggregation weights based on three similarity metrics and the reconstruction similarity associated with each client. By utilizing experiences sampled from the simulated environment, the server engages in learning a collaborative defense strategy aimed at minimizing empirical loss. This learning process employs the RL algorithm TD3 (Fujimoto, Hoof, and Meger 2018).
@@
Our Krum Median C-Median FLtrust Clipping
MNIST 1.652 1.744 1.069 1.556 1.107 1.327
Cifar10 7.201 8.698 3.537 4.469 3.589 3.696
@@
Computational Complexity
Compared to classical FL, AdaAggRL’s increased computational complexity mainly stems from $\mathrm{EnvironmentalCues}$ step. In IG algorithm, computation involves model forward propagation, loss function calculation, and backpropagation for optimization. The computational complexity depends on model complexity $M$, optimization rounds $max\_iters$, and reconstructed images $num\_images$, totaling $O(max\_iters\times M\times num\_images)$. Extracting image features via pre-trained CNN incurs a complexity of $O(C_{CNN}\times num\_images)$. Considering MMD’s constant complexity $O(C_{MMD})$, the overall computational complexity for $\mathrm{EnvironmentalCues}$ is $O(|C^{t}|\times num\_images\cdot(M\times max\_iters+C_{CNN}))+O(3|C^{t}|\cdot C_{MMD})+O(1)$.
@@
Experiments
(a) IPM
(b) LMP
(c) EB
(d) RL-attack
Figure 4: The testing accuracy variation of the global model on Cifar10 dataset under four attacks.
(a) LMP
(b) EB
(c) LMP
(d) EB
Figure 5: The testing accuracy of FL methods on MNIST-0.5 under LMP and EB as the proportion of malicious clients increases (a-b). The defense performance of AdaAggRL on MNIST-0.5 compared to the case where $S_{cl}$ is not considered under LMP (c) and the case where $S_{cg}$ and $S_{lg}$ are not considered under EB (d).
(a) MNIST-0.1, IPM
(b) MNIST-0.1, LMP
(c) MNIST-0.1, EB
(d) MNIST-0.1, RL-attack
(e) MNIST-0.5, IPM
(f) MNIST-0.5, LMP
(g) MNIST-0.5, EB
(h) MNIST-0.5, RL-attack
Figure 6: The performance of FL defense algorithms under different distribution conditions.
Experimental Settings
Dataset
@@
We conduct experiments on four datasets: MNIST (LeCun et al. 1998), F-MNIST (Xiao, Rasul, and Vollgraf 2017), EMNIST (Cohen et al. 2017), and Cifar10 (Krizhevsky, Hinton et al. 2009). Addressing the non-i.i.d. challenge in FL, we follow the approach from prior work (Fang et al. 2020) by distributing training examples across all clients. Given an M-class dataset, clients are randomly divided into M groups. The probability $q$ of assigning a training sample with label $l$ to its respective group is set, with the probability of assigning it to other groups being $\frac{1-q}{M-1}$. Training samples within the same client group adhere to the same distribution. When $q=1/M$, the distribution of training samples across M groups is uniform, ensuring that all clients’ datasets follow the same distribution. In cases where $q>1/M$, the datasets among clients are not identically distributed. Using MNIST dataset, we set $q=0.5$ to distribute training samples among clients unevenly, denoted as MNIST-0.5. MNIST-0.1 represents the scenario where MNIST is evenly distributed among clients ($q=0.1$).
Metrics
@@
We assess FL defense methods by evaluating the global model’s image classification accuracy after 500 epochs of FL training since these attacks aim to diminish testing accuracy. Higher accuracy of the global model under various attacks indicates stronger defense robustness.
Baselines
@@
To verify the effectiveness and stability of AdaAggRL, we mainly compare it with five other defense algorithms: Krum (Blanchard et al. 2017), coordinate-wise median (Median) (Yin et al. 2018), norm clipping (Clipping) (Sun et al. 2019), an extension of the vanilla coordinate-wise median (C-Median) where a norm clipping step is applied (Li, Sun, and Zheng 2022), and FLtrust (Cao et al. 2020). Krum filters malicious updates at the client level, Clipping performs gradient clipping on parameter updates before aggregation, Median and C-Median select the median or clipped median of individual parameter values from all local model updates as global model parameters, and FLtrust requires the server to have access to an amount of root data. To further illustrate its performance, AdaAggRL is compared with Feddefender (Park et al. 2023) and FedVal (Valadi et al. 2023) on CIFAR-10 dataset.